OpenStudy (anonymous):

1. Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x). x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 3 4 5 6 @uri

OpenStudy (anonymous):

2.5x-10.5 and the other side of the table is Gx=64(.5^x)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@uri @TheAsker2002 @pappi.mac @mathmate @sunshine201

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

somone help me!! im so confused

OpenStudy (anonymous):

Plug the numbers in where you see x. For example: f(2) = 2.5*2 - 10.5 = -5.5 g(2) = 64(0.5*2) = 64 f(3) = 2.5*3 - 10.5 = -3 g(3) = 64(0.5*3) = 96 Just continue the process for 4, 5 and 6.

OpenStudy (anonymous):

oh okay thanks!!

OpenStudy (anonymous):

so wait do i put the equation there or the answer?

OpenStudy (anonymous):

The answer you get after plugging in the numbers is what you put in the graph.

OpenStudy (anonymous):

there is no graph its a table.

OpenStudy (anonymous):

Table I meant

OpenStudy (anonymous):

how did you get 64 i got 16 for g2

OpenStudy (anonymous):

@sunshine201

OpenStudy (anonymous):

64(0.5 * 2) = 64(1) = 64

OpenStudy (anonymous):

64(0.5*3) = 64 (1.5) = 96 Do you get it @knightmare6?

OpenStudy (anonymous):

but its .5 not 1.5?@sunshine201

OpenStudy (anonymous):

okay but can you help me on my other question i got it now if its not 1.5 then its decaying and its not

OpenStudy (anonymous):

You have to multiply by the same number on one side that you did on the other. So because we multiplied by two in one equation, we have to do it again. We do it anywhere we see x.

OpenStudy (anonymous):

but its G(x)=64(.5^x not .5x

OpenStudy (anonymous):

Oh! Well then yes it is 16 for g2. I thought it was 0.5x. You should be more clear about how you ask your questions. There is a big difference between the two.